OpenStudy (anonymous):

can someone help me with this 4 part word problem. 1. Imagine you are in the market for a new home and are interested in a new housing community under construction in a different city. a. The sales representative informs you that there are 56 houses for sale with two floor plans still available. Use x to represent floor plan one and y to represent floor plan two. Write an equation that illustrates the situation. b. The sales representative later indicates that there are three times as many homes available with the second floor plan than the first. Write an equation that illustrates this sit

OpenStudy (anonymous):

x+y=56

OpenStudy (anonymous):

y=3x hence x+y=56 4x=56 x =14, y=42

OpenStudy (anonymous):

Him1618 what answer goes with what Please verify I'm really lose on this problem

OpenStudy (anonymous):

X + y = 56 3x = y So 4x = 56 x = 14 y = 42 therefore, there are 14 house with floor plan 1 and 42 houses with floor plan 2

OpenStudy (anonymous):

The equations are x + y = 56 and y = 3x